OpenStudy (anonymous):

solve for s s(s+4)=96

OpenStudy (anonymous):

s^2+4s-96=0 (s+12)(s-8)=0 s=-12 or 8

OpenStudy (anonymous):

s^2+4s-96=0 ===> (s+12)(s-8)=0 ==> s=-12 or s=8